# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v1

@@ VI-B Model introspection
The uncertainties obtained from probabilistic models can provide an insight into the reliability of a model when the correlation between uncertainty and model accuracy is strong. This relationship has been formalized by Mukhoti et al. [149] through two conditional likelihoods. Firstly, the accuracy given a certain prediction, $p(\mathrm{A}|\mathrm{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\mathrm{U}|\mathrm{I})$. Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{ac}$, $u_{au}$, $u_{ic}$, $u_{iu}$, respectively. Consequently, the authors combine $p(\mathrm{A}|\mathrm{C})=\frac{n_{ac}}{n_{ac}+n_{au}}$ and $p(\mathrm{U}|\mathrm{I})=\frac{n_{ic}}{n_{ic}+n_{iu}}$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
$\mathrm{PAvPU}=\frac{n_{ac}+n_{au}}{n_{ac}+n_{ic}+n_{ic}+n_{iu}}.$ (38)
This downstream task has been applied to (video) street scene [150, 151], remote sensing [152, 153], instance segmentation of various objects [154], point-cloud indoor scenes [155] brain [156, 157, 158, 17, 159, 160, 161], Multiple Sclerosis [162], cardiac [163, 164], heart ventricle [161], prostate [161], carotid artery [165], lumbosacral [19] MRI, and Optical Coherence Tomography [166, 167], skin imaging [168, 169], lung [169] and liver CT [170] and MRI  [171], and ultrasound [124, 172]. Concrete dropout has been applied to instance segmentation of C. Elegans assays [173] and street-scenes [149]. To a lesser degree, ensembling [174, 124], Variational Inference both in 2D [175, 176] and 3D [177], M-heads (auxiliary networks) [14, 178, 179], and test-time augmentation [124] have also been used to quantify the uncertainty for quality assessment.
It can be noted that uncertainty is usually only obtained on a pixel basis, while crucial information can be present in structural statistics. Hence, the Coefficient of Variation (CV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples. Also, Roy et al. [160] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to then determine the pair-wise average overlap between all respective samples as
$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\left\{\mathrm{Dice}(\mathbf{Y}_{i}=t(\mathbf{y}),\mathbf{Y}_{j}=t(\mathbf{y}))\right\}_{i\neq j}\right].$ (39)
@@ VI-C Model generalization
As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance due to the model combining effect. For instance, dropout layers at the deeper decomposition levels of the SegNet [4] improves model performance [150]. Literature also reports improved performance with Concrete Dropout for semantic (instance) segmentation [149, 173]. This has been observed in multiple domains including out-/indoor scene understanding [150, 149, 184, 185, 179], brain tumor MRI [157, 186, 159, 175], Optical Coherence Tomography [166], low dose Computed Tomography of lung nodules [187], colectoral polyps [188], cardiac MRI [18] and C.elegans roundworm microscopy images [173]. Ng et al. [18] benchmark multiple techniques for cardiac MRI segmentation and find that ensembling results in best performance improvement, while Bayes By Backprop [8] is more robust to noise distortions.
Furthermore, the improved generalization from ensembling has also shown to produce more calibrated outputs [161]. In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits [49]. Nonetheless, individual models in conventional ensembles receive data in an unstructured manner. However, it is also possible to assign specific subsets of the data to particular models (so-called ‘experts’) in the ensemble [127], commonly referred to as a Mixture of Experts (MoE). While MoE resembles ensembling in many ways, the approach additionally relies on a learnable gate that inherits the decision-taking logic. Pavlitskaya et al. [179] show merits of such approach has been observed in urban outdoor scene segmentation. For the optic cup segmentation, Ji et al. [135] also condition on a normalized expertness vector, where each each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net. Gao et al. [78] introduced Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach. Nevertheless, the methodology addresses aleatoric uncertainty quantification and has hence been discussed in Section IV-B2.
@@ VI-D Active Learning
can provide a notion to describe information gain, and thus the uncertainty, from specific datapoints. Notably, minimizing the expected increase in posterior entropy is equivalent to maximizing the mutual information between the data and model parameters. This enables reformulation in terms of the output space, rather than the complex parameter space and is known as Bayesian Active Learning by Disagreement (BALD) [191], denoted as
@@
$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]].$ (41a)
Active Learning through uncertainty quantification has been been beneficial for medical [192, 193, 194, 195, 196, 197, 198, 199] multi-view [200], remote sensing [201], street-views [202] and 3D point-cloud data [203]. These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [183] and novel temporal-ensembling methods [204] have also been used to reduce required human labeling. Some methods extend beyond pixel-level uncertainty, incorporating boundary information in the uncertainty quantification process [202, 198]. For example, Kasarla et al. [202] make use of both image and pixel-level uncertainty. Furthermore, pixels are weighed by their closeness to edges, as boundary cases are more likely to be uncertain. Similarly, Ma et al. [198] combine target and boundary-based uncertainty sampling, ensuring diversity, effective and balanced utilization of the available information.
@@ VII Discussion
This section builds on the preceding, by highlighting the main trends and discussion points in the field. The challenges and limitations related to the methods (Section VII-A) and downstream applications (Section VII-B) are discussed. Based on these observations, recommendations for future work will be provided in Section VII-C.
VII-A Methods
The core challenge with modeling segmentation distributions entails calibrated uncertainty estimates and secondly, producing spatial coherence in its samples. While the former can be achieved with conventional CNN models, the latter requires modeling correlation across pixels of the segmentation mask. As discussed in earlier sections, the desired nature of the uncertainty estimates governs the required methodology. Hence, our discussion will follow a similar format as the preceding sections, i.e. the aleatoric and epistemic approaches will initially be treated in a separate manner.
@@ VII-A1 Aleatoric methods
The literature overview indicates that three distinct routes can be taken to model the correlation. The first method is to model the correlation directly in pixels space. The second method entails using latent-variable modeling, where the correlation is encoded in latent space. Finally, test-time augmentation is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models. We will discuss the strengths and weaknesses of each of those approaches. A summary of this is presented in Table II.
@@
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification. Nonetheless, the shortcomings of VAEs are well-known. For example, such models suffer from inference suboptimality related to ELBO optimization [205, 59] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [65, 79, 70], which is hypothesized to be caused by excessively strong decoders [80]. This is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [71], residual connections [71, 72] or deep supervision [72] are required for generalization. A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotator styles or the exploration of low-likelihood regions. Hence, VAE-based models serve as good choice for the task given its shortcomings are sufficiently addressed.
Compared to the VAE-based models, the adoption of DDPMs is rather limited. This is regardless of the fact that it outperforms the VAE-based methods. Besides the point that DDPMs emerged much later in the field, its crucial limitation is regarding the tedious sequential inference procedure [92, 206, 207]. This shortcoming is exacerbated in supervised settings, which often validate through sampling on a separate data split. Furthermore, it can be noted that the best performing DDPM models are discrete in nature. While it is debatable whether shifting to categorical distributions is required for complex image generation [134], this observation does signals further investigation of the merits of categorical distributions in segmentation settings. This can visually be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13). Regardless, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g. blurry reconstructions of VAEs).
@@
Implicit methods such as test-time augmentation (TTA) [102, 103, 104, 105] have received some use in literature but have quickly been surpassed by alternative methods. However, TTA does lead in the category of simplicity, requiring almost no additional mechanisms or modifications to the employed architecture. Furthermore, CAR [56] is a lesser used method and this is likely due to well-known training problems of GAN-based models [209]. Similar to inference suboptimality in VAEs, GANs similarly required additional heuristic terms in the training objective due to instability. For example, the training objective of CAR is a summation of four separate losses.
SSNs [50] directly model correlations in pixel space and deems a simple, fast and model-agnostic modeling. Notably, SSNs also suffer from training instability due to the invertibility requirement of the covariance matrix. This can often simply be circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular. In practice, nevertheless, this quick-fix solution was much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
@@ VII-A2 Epistemic methods
In the context of segmentation, almost all discussed literature employ approximations of Variational Inference. In fact, usage of MC Dropout dominates the realm of epistemic uncertainty quantification, mainly due it being a relatively simple, cheap and straightforward approach. Nevertheless, MC dropout has been subject to substantial criticism [210, 211, 108, 122]. For example, it has been shown that MC dropout can assign 0 probability to the true posterior, which can also erroneously possesses multi-modality [210]. Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate rather than the observed data [211, 108]. These weaknesses of dropout have been used as a basis for alternative dropout techniques where the dropout rate is learned [108, 122]. Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of a functional model. If accurate uncertainty quantification is critical, MC dropout should be avoided altogether.
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling [187, 160], while in other works ensembling excels [161]. Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout when evaluating with the PAvPU metric [149]. All things considered, the preference for a particular methodology is seems to carry a strong data-dependency [124, 14]. Our recommendation is to experiment with both explicit VI and approximations such as MC Dropout and ensembling.
VII-B Applications
Up until now, theoretical insights have mostly been discussed. In this section, a deeper dive is taken into the domain-/downstream- level applications of uncertainty quantification. See Table III for an overview. Notably, most models are employed in healthcare use cases, where ambiguity frequently arises due to the trade-off between accuracy and the incisiveness of medical diagnosis systems. Furthermore, in the automotive industry, sensors often operate under constraints, and objects of interest are typically at a considerable distance which induces ambiguity in the acquired images. It is evident that in these fields, uncertainty is primarily utilized to quantify observer variability and/or to correlate uncertainty with prediction accuracy. Research on improved generalization and uncertainty-based active learning is much sparser. In the following sections, each respective downstream task is further discussed.
@@ VII-B Applications
Observer variability Model introspection Active Learning Model generalization
Lung [50, 58, 64, 65, 66, 68, 70, 71, 72, 136, 137, 76, 138, 139, 93] [169, 175, 187] [180] [187]
Brain [50, 97, 95, 140, 93, 98, 137, 142] [161, 157, 159, 158, 175] [195] [157, 186, 159, 175]
Outdoor scenes [71, 50, 145, 100, 154] [150, 151, 155, 149, 179, 152, 153] [200, 202] [149, 179, 150, 184, 185]
Cardiovascular [163, 164, 174, 161, 176, 124, 165] [192] [212]
Prostate [143, 73, 72, 136, 137] [161] [192]
@@ VII-B1 Observer variability
Varying delineation hypotheses present themselves as “noise” in the ground-truth of the data. This observer variability is often a result of viewing limitations and therefore heavily correlate to the input image. In turn, this explains the success of learning input-conditional observer variability directly through explicitly modeling the likelihood distribution, rather than to infer the underlying latent parameter distribution. Nonetheless, this approach does have its limitations. For example, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles [213].
A significant challenge in this domain is the lack of standardization, making benchmarking extremely difficult. The reasoning for this is twofold. Firstly, it is evident from Table III that a wide range of datasets are used in literature, often involving proprietary in-house data. This makes replicating the presented results impossible. Secondly, consensus on data splitting is also lacking on the publicly available datasets. For example, significant incongruity across literature is observed for the LIDC-IDRI dataset regarding data splitting and preprocessing. Similarly, there is a lack of agreement on the implementation for the GED and HM-IoU metric. Several factors, including the choice of kernel, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
Regarding the evaluation metrics, we find that the literature often places excessive emphasis on improving them. This essentially forces the algorithm to predict segmentations identical to the available ground-truth masks. This can be problematic, as it defeats the original goal to predict plausible unseen segmentations, becoming a textbook case of Goodhart’s law: ”When a measure becomes a target, it ceases to be a good measure”. Furthermore, the quality of the GED evaluations have been subject to substantial criticism as well [214, 71, 79]. Hence, practitioners should be extra vigilant and consider the importance to inquire with domain-level experts for qualitative evaluation. Because the GED and HM-IoU evaluations are dependent on on the number of segmentation masks available an alternative is to involve additional annotators per data point for more accurate evaluation, but is likely to be equally expensive. This fact additionally calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.

### v3

@@ VI-B Model introspection
The uncertainties obtained from probabilistic models can provide an insight into the reliability of a model when the correlation between uncertainty and model accuracy is strong. This relationship has been formalized by Mukhoti et al. [151] through two conditional likelihoods. Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$. Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{\text{ac}}$, $u_{\text{au}}$, $u_{\text{ic}}$, $u_{\text{iu}}$, respectively. Consequently, the authors combine $p(\text{A}|\text{C})=n_{\text{ac}}/(n_{\text{ac}}+n_{\text{au}})$ and $p(\text{U}|\text{I})=n_{\text{ic}}/(n_{\text{ic}}+n_{\text{iu}})$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
$\text{PAvPU}=\frac{n_{\text{ac}}+n_{\text{au}}}{n_{\text{ac}}+n_{\text{au}}+n_{\text{ic}}+n_{\text{iu}}}.$ (37)
This application has been applied in combination with MC dropout to model (video) street scenes [152, 153], remote sensing [154, 155], instance segmentation of various objects [156], point-cloud indoor scenes [157], brain [158, 159, 160, 17, 161, 162, 163], Multiple Sclerosis [164], cardiac [165, 166], heart ventricle [163], prostate [163], carotid artery [167], lumbosacral [19] MRI, and Optical Coherence Tomography [168, 169], skin imaging [170, 171], lung [171] and liver CT [172] and MRI  [173], and ultrasound [125, 174]. Concrete dropout has been applied to instance segmentation of C. elegans assays [175] and street-scenes [151]. To a lesser degree, ensembling [176, 125], Variational Inference both in 2D [177, 178] and 3D [179], M-heads (auxiliary networks) [14, 180, 181], and test-time augmentation [125] have also been used to quantify the uncertainty for quality assessment.
It can be noted that uncertainty is usually only obtained on pixel basis, while crucial information can be present in structural statistics. Hence, the Coefficient of Variation (CV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples. Also, Roy et al. [162] propose to evaluate structural uncertainties by assuming predictions to be thresholded to binary masks with some function $t$, to determine the pair-wise average overlap between all respective samples as
$\overline{\mathrm{D}}=\mathbb{E}_{p_{\mathbf{Y}|\mathbf{X}^{*}}}\left[\,\text{Dice}(t(\mathbf{y}_{i}),t(\mathbf{y}_{j}))\,\right]\quad\forall i\neq j.$ (38)
@@ VI-C Model generalization
As mentioned earlier in Section V-A, sampling new parameter permutations often improves segmentation performance, due to the model combining effect. For instance, dropout layers at the deeper decomposition levels of the SegNet [4] improves model performance [152]. Literature also reports improved performance with Concrete Dropout for semantic (instance) segmentation [151, 175]. This has been observed in multiple domains, including in-/outdoor scene understanding [152, 151, 186, 187, 181], brain tumor MRI [159, 188, 161, 177], Optical Coherence Tomography [168], low-dose Computed Tomography of lung nodules [189], colectoral polyps [190], cardiac MRI [18] and C. elegans roundworm microscopy images [175]. Ng et al. [18] have benchmarked multiple techniques for cardiac MRI segmentation and have found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions.
Furthermore, the improved generalization from ensembling has also shown to produce more calibrated outputs [163]. In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits [49]. Nonetheless, individual models in conventional ensembles receive data in an unstructured manner. However, it is also possible to assign specific subsets of the data to particular models (so-called ‘experts’) in the ensemble [128], commonly referred to as a Mixture of Experts (MoE). While MoE resembles ensembling in many ways, the approach additionally relies on a learnable gate that inherits the decision-taking logic. Pavlitskaya et al. [181] show merits of such approach for urban outdoor scene segmentation. For optic cup segmentation, Ji et al. [137] also condition on a normalized expertness vector, where each element corresponds to the weight given to a particular expert, and is inserted at the deepest layer of a U-Net. Gao et al. [81] have introduced a Mixture of Stochastic Experts (MoSe), which can be regarded as a stochastic adaptation of the MoE approach. Nevertheless, the methodology addresses aleatoric uncertainty quantification and has hence been discussed in Section IV-B2.
@@ VI-D Active Learning
can provide a notion to describe information gain, and thus the decrease of uncertainty for specific datapoints. Notably, minimizing the expected increase in posterior entropy is equivalent to maximizing the mutual information between the data and model parameters. This enables reformulation in terms of the output space, rather than the complex parameter space and is known as Bayesian Active Learning by Disagreement (BALD) [193], denoted as
@@
$=H[\mathbf{y}|\mathbf{x}^{*},\mathcal{D}]-\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,].$ (41)
Active Learning through uncertainty quantification has been been beneficial for medical [194, 195, 196, 197, 198, 199, 200, 201], multi-view [202], remote sensing [203], street-views [204] and 3D point-cloud data [205]. These works mostly utilize MC Dropout to obtain a notion of uncertainty, but explicit VI together with BALD [185] and novel temporal-ensembling methods [206] have also been used to reduce the required human labeling. Some methods extend beyond pixel-level uncertainty, incorporating boundary information in the uncertainty quantification process [204, 200]. For example, Kasarla et al. [204] make use of both image and pixel-level uncertainty. Furthermore, pixels are weighed by their closeness to edges, as boundary cases are more likely to be uncertain. Similarly, Ma et al. [200] combine target and boundary-based uncertainty sampling, thereby ensuring diversity, effective and balanced utilization of the available information.
@@ VII Discussion
This section highlights the main trends and discussion points in the field. The challenges and limitations related to the methods (Section VII-A) and downstream applications (Section VII-B) are discussed. Based on these observations, recommendations for future work will be provided in Section VII-C.
VII-A Discussion on methods
The core challenges with modeling segmentation distributions entails obtaining calibrated uncertainty estimates and secondly, designing spatial coherence between pixels of the samples. While the former can be achieved with conventional CNN models, the latter requires modeling correlation across pixels of the segmentation mask. As discussed in earlier sections, the desired nature of the uncertainty estimates governs the required methodology. Similarly, our discussion will follow a comparable format as the preceding sections, i.e. the aleatoric and epistemic approaches will initially be treated in a separate manner.
@@ VII-A1 Aleatoric methods
The literature overview indicates that three distinct routes can be taken to model the correlation. The first method is to model the correlation directly in the pixel space. The second method entails using latent-variable modeling, where the correlation is encoded in the latent space. Finally, TTA is a more straightforward and practical approach to obtain uncertainty and requires the least modifications to existing models. We will discuss the strengths and weaknesses of each of those approaches. A summary of this is presented in Table II.
@@
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification. Nonetheless, the shortcomings of VAEs are well known. For example, such models suffer from inference suboptimality related to ELBO optimization [207, 65] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [73, 63, 68]. This has been hypothesized to be caused by excessively strong decoders [64] and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [74], residual connections [74, 75], or deep supervision [75] are required for generalization. A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions. Hence, VAE-based models serve as good choice for the task given their shortcomings are sufficiently addressed and partially solved.
Compared to the VAE-based models, the adoption of DDPMs is rather limited. This is regardless of the fact that it outperforms the VAE-based methods. Besides the point that DDPMs emerged much later in the field, their crucial limitation, similar to thee PixelCNN [53], is the tedious sequential inference procedure [92, 208, 209]. This shortcoming is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Regardless of these shortcomings, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g. blurry reconstructions of VAEs). Furthermore, it can be noted that the best-performing DDPM models are discrete in nature. While it is debatable whether shifting to categorical distributions is required for complex image generation [136], this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings. Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
@@
Implicit methods such as TTA [103, 104, 105, 106] have received some use in literature, but have been quickly surpassed by alternative methods. However, TTA does lead in the category of simplicity, requiring no additional mechanisms or modifications to the employed architecture. Furthermore, CAR [56] is a lesser used method and this is likely due to well-known training problems of GAN-based models [211]. Due to this, GANs require additional heuristic terms in the training objective due to instability. For example, the training objective of CAR is a summation of four separate losses.
SSNs [50] directly model correlations in pixel space and deem simple, fast and model-agnostic modeling. Notably, SSNs also suffer from training instability due to the invertibility requirement of the covariance matrix. This can be often simply circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular. In practice, nevertheless, this quick-fix solution is much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
@@ VII-A2 Epistemic methods
Almost all discussed literature quantifying epistemic uncertainty employ approximations of VI. In fact, usage of MC Dropout dominates, mainly due it being a relatively simple, inexpensive and straightforward approach. Nevertheless, MC dropout has been subject to substantial criticism [212, 213, 109, 123]. For example, it has been shown that MC dropout can assign a probability of zero to the true posterior, which can also erroneously possess multi-modality [212]. Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate, rather than the observed data [213, 109]. These weaknesses have been used as a basis for alternative techniques, where the dropout rate is learned [109, 123]. Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of an already functional model. If accurate uncertainty quantification is critical, MC dropout should be avoided in any case.
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling [189, 162], while in other works ensembling excels [163]. Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout when evaluating with the PAvPU metric [151]. All things considered, the preference for a particular methodology seems to carry a strong data-dependency [125, 14]. Our recommendation is to experiment with both explicit VI and its approximations.
VII-B Discussion on applications
Up until now, theoretical insights have mostly been discussed. In this section, a deeper dive is taken into the applications of uncertainty quantification. Table III provides an overview of these applications in various domains. Notably, most models are employed in healthcare use cases, where ambiguity frequently arises due to the trade-off between accuracy and the incisiveness of medical diagnosis systems. Furthermore, in the automotive industry, sensors often operate under constraints, and objects of interest are typically at a considerable distance which induces ambiguity in the acquired images. It is evident that in these fields, uncertainty is primarily utilized to quantify observer variability and/or to correlate uncertainty with prediction accuracy. Research on improved generalization and uncertainty-based active learning is much sparser. In the following sections, each respective downstream task is further discussed.
@@ VII-B Discussion on applications
Observer Model Active Model
variability Introspection Learning generalization
Lung [50, 69, 59, 73, 60] [171, 177, 189] [182] [189]
[62, 68, 74, 75, 138]
[140, 141, 94, 139, 79]
Brain [50, 98, 96, 142] [163, 159, 161, 160, 177] [197] [159, 188]
[99, 139, 144, 94] [161, 177]
Outdoor scenes [74, 50, 147] [152, 153, 157, 151, 181] [202, 204] [151, 181, 152]
[101, 156] [186, 187]
Cardiovascular [165, 166, 176, 163] [194] [214]
[125, 167, 178]
Prostate [145, 76, 75] [163] [194]
Prostate [138, 139]
@@ VII-B1 Observer variability
Varying delineation hypotheses present themselves as “noise” in the ground-truth of the data. This observer variability is often a result of viewing limitations and therefore heavily correlate to the input image. In turn, this explains the success of directly learning input-conditional observer variability, through explicitly modeling the likelihood distribution, rather than inferring the underlying latent parameter distribution. Nonetheless, this approach does have its limitations. For example, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles [215].
A significant challenge in this domain is the lack of standardization, making benchmarking extremely difficult. The reasoning for this is twofold. Firstly, it is evident from Table III that a wide range of datasets are used in literature, often involving proprietary in-house data. This inherently makes replicating the presented results impossible. Secondly, consensus on data processing is also lacking on the publicly available datasets. For example, significant incongruity across literature is observed for the LIDC-IDRI dataset, regarding data splitting and preprocessing. Similarly, there is a lack of agreement on the implementation for the GED and HM metric. Several factors, including the kernel choice, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
Regarding the evaluation metrics, we find that the literature often places excessive emphasis on improving them. This essentially forces the algorithm to predict segmentations identical to the available ground-truth masks. This can be problematic, as it defeats the original goal to predict plausible unseen segmentations, becoming a textbook case of Goodhart’s law: ”When a measure becomes a target, it ceases to be a good measure”. Furthermore, the quality of the GED evaluations have been subject to substantial criticism as well [216, 74, 63]. Hence, practitioners should be extra vigilant and consider the importance to inquire domain-level experts for qualitative evaluation. Because the GED and HM are dependent on on the number of segmentation masks available, an alternative is to involve additional annotators per datapoint for more accurate evaluation, but is likely to be equally expensive. This aspect calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.